Ask your own question, for FREE!
Mathematics 13 Online
OpenStudy (anonymous):

Find P(not a 1). 1 2 3 4 P(X) 0.30 0.40 0.15 0.15

OpenStudy (anonymous):

\[P(\text{not 1})=1-P(1)\]

OpenStudy (anonymous):

It says the answer is 0.7 how is that possible?

OpenStudy (anonymous):

because you either get a one, or you do not if the probability you get a 1 is 0.3 then the probability you do not get a 1 is \(1-.3=.7\)

OpenStudy (anonymous):

Okay! thanks alot <3

OpenStudy (anonymous):

think of it this way if it rains 30% of the time, then it does not rain 70% of the time make sense right? yw

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

good! are there more?

OpenStudy (anonymous):

I have one more

Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!
Latest Questions
Arriyanalol: i need help
1 day ago 7 Replies 1 Medal
Arriyanalol: CONGRATS to me
1 day ago 8 Replies 2 Medals
Legomaster346: What colleges are good that are near Arab Alabama?
1 week ago 5 Replies 0 Medals
Arriyanalol: i need help
1 week ago 6 Replies 2 Medals
Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!